the age of two friends A&B differs by 3 yrs.A's father D is twice as old as A and B is twice as old as sister C. the ages of C & D differ by 30 yrs. Find their ages of A & B?

Answer by god2012(113)

Solution:
Given : A - B = 3 ------------eq.1
D = 2A ------------eq.2
B = 2C or C = B/2 ----eq.3
D - C = 30 -----------eq.4
Put eq.2 , eq.3 in eq.4, we get
2A - B/2 = 30
4A - B = 60 ------------------eq.5
Solving eq.1 and 5, we get,
A = 19
B = 16